OpenStudy (anonymous):

What is 0.35 written as a fraction in simplest form?

OpenStudy (destinymasha):

7/20

OpenStudy (destinymasha):

step 1 ] write as a fraction example --- 0.35/1

OpenStudy (destinymasha):

step 2 ] multiply the numerator and denominator by 100 so , 0.35/1 x 100 = 35/100

OpenStudy (destinymasha):

step 3 ] Simplify --- 35/100 35 divided by 5 = 7 , and 100 divided by 5 = 20

OpenStudy (destinymasha):

step 4 ] write as a fraction -- 7/20

OpenStudy (destinymasha):

Thanks for the medal!